OpenStudy (anonymous):

I Am Awarding Medals! Solve for x in the proportion x over the quantity x plus 12 equals 1 over the quantity x minus 1.

OpenStudy (anonymous):

The answer choices are x ≈ −6.02 and x ≈ 5.02 x ≈ −3.46 and x ≈ 3.46 x ≈ −3.14 and x ≈ 4.14 x ≈ −2.61 and x ≈ 4.61 I'm pretty sure you're the best one to help @satellite73 , you and your nice green bike are the best

OpenStudy (anonymous):

\[\frac{x}{x+12}=\frac{1}{x-1}\]?

OpenStudy (anonymous):

Thank you you're finally here!

OpenStudy (anonymous):

is that the question?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

I also posted the answer choices

OpenStudy (anonymous):

kk cross multiply as a first step

OpenStudy (anonymous):

\[x^2-x=x+12\] then solve the quadratic equation \[x^2-2x-12=0\]

OpenStudy (anonymous):

you should get \(x=1\pm\sqrt{13}\) and then a calcuator

OpenStudy (anonymous):

looks like about \(-2.6\) and \(4.6\)

OpenStudy (anonymous):

last ones i guess

OpenStudy (anonymous):

Thank you soooo much ! @satellite73